# Domain coarsening in fractonic systems: a cascade of critical exponents

### v1

@@ III Domain growth exponents in presence of multipole conservation
In keeping with these general considerations, we derive the domain coarsening exponents using a minimal model of multipole-conserving dynamics with a scalar order parameter. First, in Sec. III.1 we review how multipole symmetry slows the transport of the order parameter. Based on this transport, we provide a heuristic argument in a 1D toy model (Sec. III.2) and develop a continuum model for $d>2$ (Sec. III.3). Using the continuum model, we extract the domain growth law both with a scaling argument (Sec. III.4) and based on the relaxation of a perturbed domain wall (Sec. III.5). We also compute the early-time corrections to the asymptotic scaling in Sec. III.6, as these become particularly severe in multipole-conserving models and are needed to interpret numerical results.
@@ III.1 Subdiffusion of the order parameter
In a system that conserves both the total U(1)(1) charge $\int\phi(x)d^{d}x$ and its dipole moment, $P^{(1)}_{i}=\int x_{i}\phi(x)d^{d}x$, the continuity equation still holds. Microscopically, however, individual charges cannot move because of the dipole constraint. Charge can only evolve through the hopping of dipoles, so the charge current must be expressed in terms of a more fundamental dipole current. The two currents are related by
@@ III.3 Continuum model
where an overall constant of proportionality has been absorbed into the time scale. The TDLG equation, or model A in the Hohenberg-Wagner nomenclature Hohenberg and Halperin (1977), represents a dissipative evolution of the order parameter towards a minimum in the free energy. [^2] At finite temperature, the dynamics should be supplemented by a noise term to model thermal fluctuations. However, as temperature is an irrelevant parameter (in the RG sense) for coarsening, one can ignore this detail Bray (1994).
@@ IV.1 Numerical methods
We focus on the ferromagnetic Ising model on a two-dimensional square lattice of linear size $L$, with the Hamiltonian given by Eq. (1). The dynamical update rules of our Monte Carlo algorithm respect locality and the dipole (or multipole) constraint. To conserve the $m$-th multipole moment, an update must exchange two identically shaped $m$-poles. Each update therefore requires an initial site and $m+1$ vectors: $m$ vectors to specify the shape of the $m$-poles that are exchanged, and one for their separation. We use $\sigma^{+}$ ($\sigma^{-}$) to represent an update flipping a down (up) spin to an up (down) spin. [^4] This notation is borrowed from the Pauli matrices commonly used in quantum spin-1/2 systems. In this notation, a generic dipole-conserving update can be written as
@@
where $\bm{r_{0}}$ is the position of the update, $\bm{r_{1}}$ specifies the size of the dipoles involved, and $\bm{r_{2}}$ specifies their separation. Examples of dipole-conserving updates are shown in Fig. 3. Locality is enforced by taking vectors of magnitude no larger than two lattice units for dipole updates, or four lattice units for quadrupole updates. Furthermore, the periodic boundary conditions imply that multipole moments are only conserved modulo $L$.
@@ VI Conclusion
In this work, we studied coarsening in spin systems with fractonic mobility constraints. Following a quench into the ordered phase of the Ising model, we tracked how the growth of domains is modified when higher moments of the order parameter are conserved, such as its dipole or quadrupole moment. Though we focused on spin systems, our results hold for conserved multipoles of any scalar order parameter. It is well known that the linear size of domains grows as $R(t)\sim t^{1/2}$ when the order parameter is not conserved (Glauber dynamics), and $R(t)\sim t^{1/3}$ when the order parameter is conserved (Kawasaki dynamics). By generalizing this theory to the multipole case, we showed that if the $m$-th moment of the order parameter is conserved, then domains grow as $R(t)\sim t^{1/(2m+3)}$.
@@
Furthermore, we provided a simple combinatorial argument that, under mild assumptions, shows that states with extensively large domains are dynamically accessible. While domains scale with system size at asymptotically long times, growth may be problematic if the thermodynamic limit is taken first, since the relevant time scales might diverge. This opens the possibility of ergodicity-breaking and aging phenomena in such systems at finite temperature, a natural direction for future investigation. Indeed, previous work on coarsening with kinetic constraints uncovered similar glassy phenomena Sollich and Evans (1999, 2003).
This work opens other natural directions for further research, including the study of coarsening in systems with modulated symmetries, the effect of long-range updates on the dynamics, and the role of defects that weakly break the conservation laws. Among the most interesting directions is the study of coarsening in multipole-conserving quantum systems Aron et al. (2009); Chandran et al. (2013); Maraga et al. (2015); Samajdar and Huse (2024); Balducci et al. (2025) as recent experiments have shown the capability of quantum simulators to access the coarsening regime in isolated systems undergoing unitary dynamics Andersen et al. (2025); Manovitz et al. (2025). By quenching or ramping the system from a disordered to an ordered phase, one can seed the phase ordering kinetics in quantum simulators and measure the dynamical exponent $z$ via imaging. Moreover, dipole conservation can be realized as an emergent symmetry in cold-atom systems subject to a tilted external potential Guardado-Sanchez et al. (2020); Scherg et al. (2021); Kohlert et al. (2023). Whether the interplay of quantum fluctuations and constraints would lead to localization is a difficult and deep research question that deserves to be addressed.

### v2

@@ III Domain growth exponents in presence of multipole conservation
In keeping with these general considerations, we derive the domain coarsening exponents using a minimal model of multipole-conserving dynamics with a scalar order parameter. First, in Sec. III.1 we review how multipole symmetry slows the transport of the order parameter. Based on this transport, we provide a heuristic argument in a 1D toy model (Sec. III.2) and develop a continuum model for $d>2$ (Sec. III.3). Using the continuum model, we extract the domain growth law both with a scaling argument (Sec. III.4) and based on the relaxation of a perturbed domain wall (Sec. III.5). We also compute the early-time corrections to the asymptotic scaling in Sec. III.6, as these become particularly severe in multipole-conserving models and are needed to interpret numerical results. Notice that we make use of a continuous time model, which inherently assumes that the final results do not depend qualitatively on the specific moves being performed in a discrete-time, numerical setting.
@@ III.1 Subdiffusion of the order parameter
In a system that conserves both the total U(1)(1) charge $\int\phi(x)d^{d}x$ and its dipole moment, $P^{(1)}_{i}=\int x_{i}\phi(x)d^{d}x$ (where we have written Eq. (8) in the continuum), the continuity equation still holds. Microscopically, however, individual charges cannot move because of the dipole constraint. Charge can only evolve through the hopping of dipoles, so the charge current must be expressed in terms of a more fundamental dipole current  [^2] In a (quantum) lattice model, the charge and dipole currents can be discretized as $J_{\text{charge}}\sim-i(\sigma^{+}_{r}\sigma^{-}_{r+1}-\text{h.c.})$ and $J_{\text{dip}}\sim-i(\sigma^{+}_{r}\sigma^{-}_{r+1}\sigma^{-}_{r+2}\sigma^{+}_{r+3}-\text{h.c.})$, where $\sigma^{\pm}$ flips a spin up/down.. The two currents are related by
@@ III.3 Continuum model
where an overall constant of proportionality has been absorbed into the time scale. The TDLG equation, or model A in the Hohenberg-Halperin-Wagner nomenclature Hohenberg and Halperin (1977), represents a dissipative evolution of the order parameter towards a minimum in the free energy. [^3] At finite temperature, the dynamics should be supplemented by a noise term to model thermal fluctuations. However, as temperature is an irrelevant parameter (in the RG sense) for coarsening, one can ignore this detail Bray (1994).
@@ IV.1 Numerical methods
We focus on the ferromagnetic Ising model on a two-dimensional square lattice of linear size $L$, with the Hamiltonian given by Eq. (1). The dynamical update rules of our Monte Carlo algorithm respect locality and the dipole (or multipole) constraint. To conserve the $m$-th multipole moment, an update must exchange two identically shaped $m$-poles. Each update therefore requires an initial site and $m+1$ vectors: $m$ vectors to specify the shape of the $m$-poles that are exchanged, and one for their separation. We use $\sigma^{+}$ ($\sigma^{-}$) to represent an update flipping a down (up) spin to an up (down) spin. [^5] This notation is borrowed from the Pauli matrices commonly used in quantum spin-1/2 systems.
A generic dipole-conserving update can be written as
@@
where $\bm{r_{0}}$ is the position of the update, $\bm{r_{1}}$ specifies the size of the dipoles involved, and $\bm{r_{2}}$ specifies their separation. This process is the analogue of the charge-conserving update $\sigma^{+}(\bm{r}_{0})\sigma^{-}(\bm{r_{0}}+\bm{r}_{1})$, which exchanges a positive and negative charge. Much like the dipole exchange of Eq. (35), a quadrupole-conserving updates corresponds to two simultaneous dipole exchanges. In this way, an $m$-pole conserving update can be hierarchically constructed from lower moments Feldmeier et al. (2020); Gliozzi et al. (2023). Examples of dipole-conserving updates are shown in Fig. 3. Locality is enforced by taking vectors of magnitude no larger than two lattice units for dipole updates, or four lattice units for quadrupole updates. Furthermore, the periodic boundary conditions imply that multipole moments are only conserved modulo $L$.
@@ VI Conclusion
In this work, we studied coarsening in spin systems with fractonic mobility constraints. Following a quench into the ordered phase of the Ising model, we tracked how the growth of domains is modified when higher moments of the order parameter are conserved, such as the dipole moment. Though we focused on spin systems, our results hold for conserved multipoles of any scalar order parameter. It is well known that the linear size of domains grows as $R(t)\sim t^{1/2}$ when the order parameter is not conserved (Glauber dynamics), and $R(t)\sim t^{1/3}$ when the order parameter is conserved (Kawasaki dynamics). By generalizing this theory to the multipole case, we showed that if the $m$-th moment of the order parameter is conserved, then domains grow as $R(t)\sim t^{1/(2m+3)}$.
@@
Furthermore, we provided a simple combinatorial argument that, under mild assumptions, shows that states with extensively large domains are dynamically accessible. While domains scale with system size at asymptotically long times, growth may be problematic if the thermodynamic limit is taken first, since the relevant time scales might diverge. This allows for ergodicity-breaking and aging phenomena in such systems at finite temperature, a natural direction for future investigation. Indeed, previous work on coarsening with kinetic constraints uncovered similar glassy phenomena Sollich and Evans (1999, 2003), which may also extend to classical fracton spin liquids Placke et al. (2024).
This work opens other natural directions for further research, including the study of coarsening in systems with modulated symmetries, the effect of long-range updates on the dynamics, and the role of defects that weakly break the conservation laws. Among the most interesting directions is the study of coarsening in multipole-conserving quantum systems Turkowski et al. (2006); Aron et al. (2009); Chandran et al. (2013); Maraga et al. (2015); Samajdar and Huse (2024); Balducci et al. (2025) as recent experiments have shown the capability of quantum simulators to access the coarsening regime in isolated systems undergoing unitary dynamics Andersen et al. (2025); Manovitz et al. (2025). By quenching or ramping the system from a disordered to an ordered phase, one can seed the phase ordering kinetics in quantum simulators and measure the dynamical exponent $z$ via imaging. Moreover, dipole conservation can be realized as an emergent symmetry in cold-atom systems subject to a tilted external potential Guardado-Sanchez et al. (2020); Scherg et al. (2021); Kohlert et al. (2023). Whether the interplay of quantum fluctuations and constraints would lead to localization is a difficult and deep research question that deserves to be addressed.